12.1 Sales forecasts.   For the prior three​ years, sales for National Beverage Company have been ​$22,059,000...

12.1

Sales forecasts.  

For the prior three​ years, sales for National Beverage Company have been

​$22,059,000 (2011), $23,186,000 (2012), and $24,032,000 (2013). The company uses the prior two​ years' average growth rate to predict the coming​ year's sales.

  1. What were the sales growth rates for 2012 and​ 2013?
  2. What is the expected sales growth rate using a​ two-year average for​ 2014? What is the sales forecast for​2014?
  3. What was the sales growth rate for​ 2012?
Answer #1

1. Sales growth rate for 2012=(sales in 2012/sales in 2011)-1=(23186000/22059000)-1=5.11%

Sales growth rate for 2013=(sales in 2013/sales in 2012)-1=(24032000/23186000)-1=3.65%

2. Two year average rate=(5.11%+3.65%)/2=4.38%

Sales forcast for 2014=2013 sales*(1+4.38%)=24032000*1.0438=25,084,357